# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 16 $\text{tempCost}\leftarrow\texttt{NewCost}(\mathbf{u,v})$;
            17 if tempCost $<\mathbf{v}.\textnormal{{cost}}$
@@ 6.2 RRT and variants for PTOPP
RRT emerges as an online and single-query counterpart of PRM since it allows for incremental sampling during planning instead of batch sampling beforehand. It maintains a tree structure and is probabilistically complete but non-optimal. RRT is the asymptotically optimal version of RRT by adding the rewiring procedure which locally modifies existing edge connections every time a new sample is inserted to ensure nodes are reached through the path of minimum cost from the root. In this way, the tree embeds optimal paths from the root for all nodes in the random tree growth process. To do this in PTOPP, the backbone of RRT is retained. Meanwhile, core modifications are introduced similarly to those in PRM for PTOPP to guarantee that the minimum cost path to each node is correctly maintained.
RRT for PTOPP is detailed in Algorithm 10 which reuses the architecture and basic routines of vanilla RRT. Again, the addition of NewCost() and SubtreeUpdate() is crucial for cost computation and dynamic tree maintenance. At first, $\mathbf{x}_{0}$ is inserted into $V$, localized in cells, and the optimal path cost $f_{c}$ is assigned to be infinity. The proposed algorithms are cost-agnostic and compatible with costs in Section 4. The specific cost is conveniently customizable in NewCost(). SubtreeUpdate() is invoked when creating the edge along the minimum cost path to the newly sampled $\mathbf{x}_{new}$ and when adjusting existing edges in the rewire routine. It alters edges and propagates resulting attribute changes to descendant nodes. Lexicographic comparison of (cost, len) is used in rewire condition (line 10 in Algorithm 10). In implementation, updating $V$ and $E$ is accomplished implicitly by updating related node attributes, i.e., parent and children. To retrieve the optimal path, only the parent of $\mathbf{x}_{g}$ needs to be tracked.
1 Input $\mathbf{x}_{0},\mathbf{x}_{g}$ and conduct passages and cell preprocessing;
@@ 9 $X_{near}\leftarrow\texttt{Near}(G=(V,E),\mathbf{x}_{new},r_{RRT^{*}})$;
            10 foreach $\mathbf{x}_{near}\in X_{near}$ do
                  11 if not ObstacleFree$(\mathbf{x}_{near},\mathbf{x}_{new})$ then
                        12 $X_{near}.\text{remove}(\mathbf{x}_{near})$;
@@
            14 $c_{min}\leftarrow\texttt{NewCost}(\mathbf{x}_{min},\mathbf{x}_{new})$;
            15 foreach $\mathbf{x}_{near}\in X_{near}$ do
                  16 if  NewCost$(\mathbf{x}_{near},\mathbf{x}_{new})<c_{min}$ then
                        17 $\mathbf{x}_{min}\leftarrow\mathbf{x}_{near}$;
                        18 $c_{min}\leftarrow\texttt{NewCost}(\mathbf{x}_{near},\mathbf{x}_{new})$;
            19SubtreeUpdate$(\mathbf{x}_{min},\mathbf{x}_{new})$;
            20 $V\leftarrow V\cup\{\mathbf{x}_{new}\};\,E\leftarrow E\cup\{(\mathbf{x}_{min},\mathbf{x}_{new})\}$;
            21 foreach $\mathbf{x}_{near}\in X_{near}$ do
@@ 20 $V\leftarrow V\cup\{\mathbf{x}_{new}\};\,E\leftarrow E\cup\{(\mathbf{x}_{min},\mathbf{x}_{new})\}$;
                        23 $E\leftarrow E\backslash\{(\mathbf{x}_{near}.\texttt{parent},\mathbf{x}_{near})\}\cup\{(\mathbf{x}_{new},\mathbf{x}_{near})\}$;
                        24 SubtreeUpdate$(\mathbf{x}_{new},\mathbf{x}_{near})$;
            25if Len$(l(\mathbf{x}_{new},\mathbf{x}_{g}))<r_{RRT^{*}}$ and ObstacleFree$(\mathbf{x}_{new},\mathbf{x}_{g})$ and $\texttt{NewCost}(\mathbf{x}_{new},\mathbf{x}_{g})<f_{c}$ then
@@ 27 $f_{c}\leftarrow\texttt{NewCost}(\mathbf{x}_{new},\mathbf{x}_{g})$;
Another asymptotically optimal algorithm closely related to RRT is the rapidly exploring random graph (RRG) which incrementally builds an undirected graph. $\mathbf{x}_{new}$ will connect to all collision-free nearby nodes $\mathbf{x}_{near}\in X_{near}$, leading to a much richer edge set that contains RRT edge set as its subset in the same sampling sequence. To adapt RRG for PTOPP, its graph construction phase remains unchanged. Subsequent optimal path search follows the workflow of the query phase in Algorithm 9. Namely, RRG for PTOPP fits into the generic pipeline in Algorithm 9 and only differs in preprocessing for graph construction. Additionally, the asymptotically optimal k-nearest variants of PRM, RRT, and RRG, in which new edges are sought to k nearest neighbors of $\mathbf{x}_{new}$ and k varies with the vertex number $|V|$, are also compatible with PTOPP and only differ from the presented algorithms in the near neighbors search procedure.
@@ 6.3.1 Probabilistic completeness
For consistency with previous literature regarding the goal configuration (Karaman and Frazzoli, 2011; Kleinbort et al., 2018), assume there is a goal region $\mathcal{X}_{goal}\subset\mathcal{X}_{free}$ and $\mathbf{x}_{g}\in\mathcal{X}_{goal}$, a path $\sigma$ reaches the goal if $\sigma(1)\in\mathcal{X}_{goal}$. Probabilistic completeness requires the probability that the algorithm finds a feasible path in any robustly feasible path planning problem, i.e., problems that allow for a feasible path with positive clearance to obstacles, converges to one as the sample number goes to infinity. It is easy to see that PRM for PTOPP, together with its variants, is probabilistic complete since the graph construction phase is not modified, and so is RRG. For avenues based on RRT, though the resulting edge set $E$ differs from the vanilla version due to new path costs, the vertex set $V$ is identical for the same sampling sequence. Meanwhile, the graph is connected by construction. Hence, its probabilistic completeness directly follows from that of RRT. Moreover, because the constructed graph is unchanged or connected, the property that the probability of failure in finding a path for a feasible planning problem decays exponentially with the sample number in the original algorithms still holds.
Figure 10: Examples of detected passage and Gabriel cell distributions in obstacle-dense environments. Passage region boundaries and passage segments are dashed segments in blue and red, respectively. (a)-(c) have similar obstacle sizes (side length is 40) and obstacle numbers are 20, 40, and 60. (d)-(f) add obstacle size variances where obstacle side lengths are integers randomized in [20, 60].
@@ 6.3.2 Asymptotic optimality
Suppose OPP problems admit a robustly optimal solution, i.e., the optimal path $\sigma^{*}$ is homotopic to some feasible path with positive clearance to obstacles, and for any sequence of feasible paths $\{\sigma_{n}\}_{n\in\mathbb{N}}$, $\lim_{n\rightarrow\infty}\sigma_{n}=\sigma^{*}$ under some well-defined path norm for path distance computation, $\lim_{n\rightarrow\infty}f_{c}(\sigma_{n})=f_{c}(\sigma^{*})$. An algorithm is asymptotically optimal if the minimum cost it finds converges to the optimal cost as the sample number approaches infinity. Since the graph construction procedure remains unchanged from the original versions in PRM, RRG, and their variants for PTOPP, the algorithms’ asymptotic optimality is inherited because path costs are compatible. More specifically, the unchanged graph structures ensure that a sequence of paths converging to the optimal path exists with probability one for all large sampling numbers. Using the monotonicity of the best path cost with sample increasing, the minimum cost these algorithms return converges to $f_{c}(\sigma^{*})$ almost surely. For PTOPP based on RRT, only the routines related to path cost computations are customized. It is guaranteed that the best path starting from $\mathbf{x}_{0}$ and reaching each new vertex $\mathbf{x}_{i}$ is found, and each vertex has a unique parent with the smallest cost. Then the asymptotical optimality follows.
@@ 6.3.3 Computational complexity
For graph construction, it has been shown that sampling-based optimal planners take $O(n\log n\log^{d}m)$ time under efficient implementations of core procedures such as collision check, nearest neighbor search, and range search. Concisely, the number of nearby nodes inside the search ball centered at the new node in iteration $n$ is of $O(\log n)$. They can be found in $O(\log n)$ time via approximate range search. Checking collisions between an edge and $m$ obstacles in $\mathbb{R}^{d}$ can be executed in $O(\log^{d}m)$ time with spatial trees. Summing up over $n$ iterations, the total computation time is of $O(n\log n\log^{d}m)$ (Karaman and Frazzoli, 2011). However, note that the complexity of the subtree update routine in RRT is not taken into account. The total edge number $|E|$ in RRT is of $O(n)$, implying that the average edge change number in the rewiring procedure is of $O(1)$. As a graph traversal routine, the complexity of SubtreeUpdate() is determined by the node number of the subtree branch. RRT maintains a tree similar to the random online nearest neighbor graph. The expected subtree node number rooted at a random node is no more than $\log n$ (Lichev and Mitsche, 2024). The total complexity thus remains.
Figure 11: Examples of detected passage and Gabriel cell distributions in 3D obstacle-dense environments. Passages are depicted as red transparent planes. (a)-(c) depict passages detected on the ground and the passage height is the lower height between obstacles constructing the passage. Obstacle numbers $m=20,40$, and 60, respectively. (d)-(f) depict passages in all height intervals.
@@
In algorithms for PTOPP, new computation overhead is primarily introduced by passage and cell detection, and passage traversal check for edges. Passage and cell detection using Delaunay graphs takes $O(m\log m)$ time. As $n\gg m$ holds in general, most new computations are induced by passage traversal check. A single check in (27) takes constant time, and the number of passages examined for each edge decides the cumulative computation time. The passage number obtained via the Gabriel condition is of $O(m)$ (Mark et al., 2008; Huang et al., 2024a). A brute-force traversal strategy yields $O(n\log n\log^{d}m+nm)$ time complexity and dramatically increases the planning time. By positioning nodes in Gabriel cells, only a constant number of passages need to be examined since the expected passage side number of a decomposed cell is of $O(1)$. Meanwhile, the average number of cells an edge passes does not exceed $O(1)$ given $n\gg m$ in $\mathbb{R}^{2}$ (note that the cell size is of $1/O(m)\times 1/O(m)$ and the edge length is of $O(\log(n)/n)^{\frac{1}{2}}$ in a planar square if the obstacle sizes are much smaller than the environment). Thereby, the time complexity remains $O(n\log n\log^{d}m)$. Summing up, the overall time complexity is $O(n\log n\log^{d}m+m\log m)$ for passage and cell detection, and graph construction in PTOPP.
@@
In the path query phase to extract the optimal path from the graph, the time complexity is the same as the shortest path problem with nonnegative edge lengths, whose asymptotic optimal bound is $O(|V|\log|V|+|E|)$ (Schrijver, 2003). Trivially, $|V|\in O(n)$ for all algorithms. $|E|\in O(n\log n)$ for PRM, RRG, and their k-nearest variants. $|E|\in O(n)$ for RRT and its k-nearest variant (Karaman and Frazzoli, 2011). Since node parent (including the goal $\mathbf{x}_{g}$) is maintained online in RRT, no separate path finding routine is required. In summary, after efficient passage detection, PTOPP is solved without increasing the time complexity of sampling-based optimal planners. As for space complexity, apart from $O(|V|+|E|)$ space for graph storage, only $O(m)$ extra space is needed to store passages and cells.
@@ 7.1 Passage detection and cell decomposition using Gabriel graphs
Table 1: Parameters in linear regression of passage and cell numbers with respect to obstacle number in different test setups. PR: Passage region, PS: Passage segment.
Test setups Pssage-PR Pssage-PS Cell-PR Cell-PS
Fixed obstacle sizes Slope 1.95 2.32 0.95 1.32
$R^{2}$ 1.00 1.00 1.00 1.00
Varying obstacle sizes Slope 1.99 2.26 0.99 1.26
$R^{2}$ 1.00 1.00 1.00 1.00
Figure 12: Passage and cell numbers detected in different planar setups. (a), (b) Passage numbers in different numbers of obstacles with a fixed side length of 40 in (a) and varying side lengths picked in $[20,60]$ in (b). (c), (d) Cell numbers with different numbers of obstacles. (e), (f) Passage and cell numbers with the same number of obstacles of different sizes.
@@ 7.1.1 Passage detection and cell decomposition results
Figure 10 showcases examples of passages and cells detected amid random obstacles in 2D maps. Passage regions are bounded by blue dashed lines. Red dashed lines are passage segments, i.e., distances, between obstacle pairs that have constructed valid passages. Two types of lines may overlap. The environment size is $1000\times 600$. In Figure 10-10, obstacles (triangles, rectangles) are of similar sizes controlled by an identical side length of 40 and are densified with an increasing obstacle number $m$. To get a complete map decomposition, environment walls are processed as obstacles in this part, and passages between adjacent walls are assumed at the four corners. In Figure 10-10, obstacles additionally vary in size with side length changes in $[20,60]$ to enrich randomness. Overall, the detected passages and cells through the proposed criterion are sparse and informative in indicating narrow regions. Passage regions depict confined locations accurately and provide more elaborate passage descriptions than simple segments. The shadow volume intersection definition in (4) endows passages with sizes insensitive to relative obstacle size differences even in the presence of significant size differences in Figure 10-10.
Spatial passages detected in 3D scenes are shown in Figure 11. Passage segments evolve into passage planes with valid height intervals between obstacles in 3D space, which resemble spatial gates. For clarity, passage region boundaries and passages incident to environment walls are not depicted. The dimension of the environment is $1000\times 6000\times 400$. Figure 11-11 plot passages only considering the obstacle distribution on the base. They start at zero height, i.e., the base, and have the upper height bound $\overline{h}$ equal to the lower obstacle forming them. This description is used in (Huang et al., 2024a). The corresponding passage distributions at all heights are shown in Figure 11-11 for comparison. Narrow passages are detected at all heights for a comprehensive description of places with restricted accessible free space. Passages traversed by paths in 3D space can thus be completely captured. Passages get sparser as height increases because fewer obstacles exist at large heights. As is observed, complex decompositions of the environment are generated, making direct node positioning challenging. This necessitates the proposed node projection localization in base cells for rapid passage traversal check during planning.
Figure 13: Passage and cell numbers detected in 3D environments. Passages and cells on the base are detected considering the planar obstacle distribution on the base for comparison.
Figure 14: Average passage detection time. (a),(b) Passage detection based on Delaunay graphs and direct traversal check. Both faster detection speeds and smaller time increases are achieved using Delaunay graphs. (c) Passage and compound cell detection in 3D space.
Table 2: Average passage and cell detection time in different test conditions (ms). Data are collected from 50 trials in each setup. Results are reported for 2D obstacles of both fixed and varying sizes. BF: Brute-force traversal passage check. DG: Using Delaunay graph. Spatial passage detection in 3D space and corresponding planar passage detection on the base are with DG.
Obstacle number 20 40 60 80 100 120 140 160 180 200
Fixed obstacle sizes BF passage detection 2.0 10.9 28.4 53.2 88.8 133.8 185.5 244.7 315.2 396.6
DG passage detection 1.9 5.8 10.0 14.5 18.9 23.5 27.8 32.3 36.7 40.7
Cell detection $<$ 0.1 0.1
Varying obstacle sizes BF passage detection 2.0 10.9 28.0 53.5 89.0 132.8 185.2 244.9 313.3 392.2
DG passage detection 1.9 5.8 10.2 14.6 19.3 23.7 28.1 32.6 37.0 41.4
Cell detection $<$ 0.1 0.4
Varying obstacle sizes (3D) Spatial passage detection 4.1 16.5 33.0 52.5 73.8 98.3 124.8 152.1 182.1 215.8
Base passage detection 2.0 6.1 10.8 15.3 20.1 24.9 29.8 34.2 38.8 43.6
Compound cell detection 4.0 12.6 24.5 38.6 55.3 75.0 89.1 118.9 137.0 154.9
@@
Statistics of passages and cells are showcased in Figure 12 where obstacle numbers are equally spaced from 20 to 200. To eliminate the influence of large environment walls, only passages between obstacles are detected. Since relative obstacle sizes are indeterminate, fixed and varying obstacle size setups are tested separately. In comparison, the detection condition that solely utilizes passage segments as a simplified passage representation is used. For each obstacle number, data are collected from 50 random obstacle distributions. The linearity between passage segments and obstacle numbers has been shown in (Huang et al., 2024a). Figure 12 and 12 demonstrate that the linear increase of passage numbers with respect to obstacle numbers also holds for the amended Gabriel condition using passage regions. Because of the linearity between triangle and edge numbers in random Delaunay graphs, the linearity between Gabriel cells and passages is expected as edges have the same probability of failing the amended Gabriel condition. As a result, the cell number is linear in the passage number and further, in the obstacle number in Figure 12, 12. Its value is roughly half the passage numbers and close to the obstacle number. In linear regression, the coefficients of determination ($R^{2}$) are larger than 0.99 in all the above relations (see Table 1), suggesting strong linearity.
Figure 15: PTOPP examples with different cost and environment setups. (a)-(c) Obstacles are uniformly distributed in the environment and obstacle numbers are 20, 40, and 60, respectively. Obstacles have varying shapes and sizes. (d)-(f) Obstacles of the same numbers respect Gaussian distributions.
@@
Next, statistics of passage detection in 3D environments are reported in Fig. 13. As the impact of relative obstacle sizes has been investigated, the general case of varying obstacle sizes is adopted. Each obstacle is assigned a random height in $[0,400]$, and the base ground remains the 2D dimension of $1000\times 600$. For comparison, the passage and cell numbers on the base ground considering the planar obstacle distribution are listed. The data show that the 3D passage number is also linear in the obstacle number. In other words, $O(m)$ valid spatial passages exist. This linearity originates from the fact that all intermediate edges in the incremental Delaunay graph construction process are checked for passages by the Gabriel condition. The ratio of the spatial passage number to the base passage number is $1.7$ on average, suggesting good sparsity of spatial passages and that the 3D passage number is not significantly larger than that in the associated planner case. Beyond the planner setups, sparse passages are distributed in 3D space via the proposed detection condition, which is crucial to enabling efficient planning in PTOPP.

### v2

@@ 11 $\text{tempCost}\leftarrow\texttt{NewCost}(\mathbf{u,v})$;
    12 if tempCost $\mathbf{v}.\textnormal{{cost}}$
@@ 6.2 RRT and variants for PTOPP
RRT emerges as the asymptotically optimal version of RRT. It introduces a rewire procedure that locally modifies existing edges every time a new node is inserted, ensuring that all nodes are reached with the minimum cost from the root. To do this in PTOPP, the backbone of RRT is retained. At first, $\mathbf{x}_{0}$ is localized in cells. Core modifications are introduced similarly to above to correctly maintain the minimum cost path to each node. Again, the addition of NewCost() and SubtreeUpdate() is crucial for cost computation and dynamic tree maintenance. The proposed algorithms are cost-agnostic and the cost type is conveniently customizable in NewCost(). SubtreeUpdate() is invoked when creating the edge along the minimum cost path to the newly sampled $\mathbf{x}_{new}$ and adjusting existing edges in the rewire routine, which alter edges and propagate attribute changes in the tree. Lexicographic comparison of (cost, len) is used in choosing the parent node and the rewire condition (line 10 in Algorithm 10). RRT for PTOPP in the architecture of vanilla RRT is detailed in Algorithm 11.
Another asymptotically optimal algorithm closely related to RRT is rapidly exploring random graph (RRG) which incrementally builds an undirected graph. $\mathbf{x}_{new}$ connects to all collision-free nearby nodes $\mathbf{x}_{near}\in X_{near}$, leading to a much richer edge set that contains RRT edge set as its subset in the same sampling sequence. To adapt RRG for PTOPP, its graph construction phase remains unchanged. Subsequent optimal path search follows the query phase in Algorithm 9. Namely, RRG for PTOPP fits into the generic pipeline in Algorithm 9 and merely differs in graph construction. Additionally, the asymptotically optimal k-nearest variants of PRM, RRT, and RRG, in which new edges are sought to k nearest neighbors of $\mathbf{x}_{new}$ and k varies with the vertex number $|V|$, are also compatible with PTOPP. They only differ in the near neighbors search procedure.
1Input $\mathbf{x}_{new}$ and its conlision-free neighbor set $X_{near}$;
2 foreach $\mathbf{x}_{near}\in X_{near}$ do
[…]
Algorithm 10 Rewire With Length-Regulation in RRT
1Input $\mathbf{x}_{0},\mathbf{x}_{g}$ and conduct passages and cell preprocessing;
@@ 9 $X_{near}\leftarrow\texttt{Near}(G=(V,E),\mathbf{x}_{new},r_{RRT^{*}})$;
    10 $\mathbf{x}_{min}\leftarrow\texttt{GetParent}(X_{near},\mathbf{x}_{new})$;
@@
    12 Rewire;
@@ 6.3.1 Probabilistic completeness
Assume there is a goal region $\mathcal{X}_{goal}\subset\mathcal{X}_{free}$, a path $\sigma$ reaches the goal if $\sigma(1)\in\mathcal{X}_{goal}$. Probabilistic completeness requires the probability that an algorithm finds a feasible path in any robustly feasible problem, i.e., problems having a feasible path of positive clearance to obstacles, converges to one as the sample number $n$ goes to infinity. It is easy to see that PRM for PTOPP, together with its variants, is probabilistic complete since the graph construction phase is not modified, and so is RRG. For RRT-based avenues, though the resulting edge set $E$ differs from the vanilla RRT due to new path costs, the vertex set $V$ is identical for the same sampling sequence. Meanwhile, the graph is connected by construction. Hence, its probabilistic completeness directly follows from that of RRT. Moreover, because the constructed graphs are unchanged or connected, the exponential decay in the failure probability of finding a path for a feasible planning problem with $n$ still holds.
@@ 6.3.2 Asymptotic optimality
Suppose an OPP problem admits a robustly optimal solution, i.e., the optimal path $\sigma^{*}$ is homotopic to some feasible path with positive clearance. An algorithm is asymptotically optimal if the minimum cost it finds converges to the optimal cost as $n$ approaches infinity. Since graph construction remains unchanged from original PRM, RRG, and their variants in PTOPP, the algorithms’ asymptotic optimality is inherited because path costs are compatible. More specifically, the unchanged graph structures ensure that a sequence of paths converging to the optimal path exists with probability one for all large $n$. Using the monotonicity of the best path cost with respect to $n$, the minimum cost these algorithms return converges to $f_{c}(\sigma^{*})$ almost surely. For PTOPP based on RRT, only path cost computations are customized. It is guaranteed that the best path starting from $\mathbf{x}_{0}$ and reaching each new vertex $\mathbf{x}_{i}$ is found and each vertex has a unique parent. Then the asymptotical optimality follows.
Figure 10: Examples of passage and Gabriel cell distributions in obstacle-dense environments. Passage region boundaries and passage segments are dashed segments in blue and red, respectively. (a)-(c) have similar obstacle sizes (side length is 40) and the obstacle number $m=$ 20, 40, and 60, respectively. (d)-(f) have obstacle side lengths randomized in [20, 60].
@@ 6.3.3 Computational complexity
Graph construction takes $O(n\log n\log^{d}m)$ time under efficient implementations of core procedures. However, the complexity of subtree update in RRT is not taken into account in the original analysis in (karaman2011sampling). The total edge number $|E|$ in RRT is of $O(n)$, implying that the average edge changes in the rewiring procedure are of $O(1)$. Doing graph traversal, the complexity of SubtreeUpdate() is determined by the node number of the subtree branch. RRT maintains a tree similar to the random online nearest neighbor graph. The expected subtree node number rooted at a random node is no more than $\log n$ (lichev2024new). The total complexity thus remains.
@@
New computational overhead in PTOPP is introduced by passage and cell detection, and passage traversal check. Passage and cell detection using Delaunay graphs takes $O(m\log m)$ time. As $n\gg m$, most new computations occur in passage traversal check. One-time check in (27) takes constant time and the number of passages examined for each edge decides the cumulative computations. The total passage number after Gabriel condition can be approximated by the edge number $O(m)$ in Gabriel graphs (mark2008computational; huang2024homotopic). Brute-force traversal yields $O(nm)$ complexity and dramatically increases the planning time. By positioning nodes in Gabriel cells, only a constant number of passages need to be examined. The average cell number an edge passes does not exceed $O(1)$ given $n\gg m$ in $\mathbb{R}^{2}$ (the cell size is $O(1/m)\times O(1/m)$ and the edge length is $O((\log(n)/n)^{\frac{1}{2}})$ in a unit square. Meanwhile, $O(1)$ passage sides exist in a cell. Thereby, the time complexity of graph construction remains. Summing up, the overall time complexity is $O(n\log n\log^{d}m+m\log m)$ for passage and cell detection, and graph construction in PTOPP.
@@
In the path query phase to extract the optimal path from the graph, the time complexity is the same as the shortest path problem with nonnegative edge lengths, whose asymptotic optimal bound is $O(|V|\log|V|+|E|)$ (schrijver2003combinatorial). Trivially, $|V|\in O(n)$ for all algorithms. $|E|\in O(n\log n)$ for PRM, RRG, and their k-nearest variants. $|E|\in O(n)$ for RRT and its k-nearest variant (karaman2011sampling). Since node parent is maintained online in RRT, no separate path finding routine is required. In summary, after efficient passage detection, PTOPP is solved without increasing the time complexity of sampling-based optimal planners. As for space complexity, apart from $O(|V|+|E|)$ space for graph storage, only $O(m)$ extra space is needed to store passages and cells.
@@ 7.1.1 Passage detection and cell decomposition results
Figure 10 showcases examples of passages and cells detected amid 2D random obstacles (map size $1000\times 600$). Passage region bounds are blue dashed lines. Red dashed lines are passage segments, i.e., distances, between obstacle pairs that construct valid passages. Two segment types may overlap. In Figure 10-10, densifying obstacles have an identical side length and thus similar sizes. To get a complete map decomposition, environment walls are processed as adjacent obstacles. In Figure 10-10, obstacles additionally vary in size to enrich randomness. Detected passages and cells are sparse and informative to depict confined locations. They provide passage descriptions more elaborate than simple segments. The proposed passage region definition in (4) also endows passages with sizes insensitive to significant obstacle size differences in Figure 10-10. Spatial passages in 3D environments featuring random obstacles are shown in Figure 11 (map size $1000\times 600\times 400$). Passage segments evolve into gates in 3D space. For clarity, passages incident to the environment walls are not depicted. Figure 11-11 plot passages only considering obstacle distribution on the base as in (huang2024homotopic). They start from the base and end at the lower incident obstacle height. In the bottom row are corresponding spatial passages that provide a comprehensive description of places with restricted accessible free space. A notable observation is that passages are sparser at large heights because fewer obstacles are there. Complex decompositions make direct node positioning hard and necessitate the proposed node projection localization in base cells.
Figure 11: Examples of passage and Gabriel cell distributions in 3D obstacle-dense environments. Passages are depicted as red transparent planes. (a)-(c) depict passages detected on the ground and the passage height equals the lower incident obstacle height. $m=20,40$, and 60, respectively. (d)-(f) depict passages in all height intervals.
@@
Statistics of passages and cells are showcased in Figure 12. Since relative sizes are indeterminate, fixed and varying obstacle sizes are tested separately with no evident statistical differences. For brevity, results in common varying-size setups are reported. In comparison, the detection condition that solely utilizes passage segments as a simplified passage representation is used. For each obstacle number $m$, data are collected from 50 random distributions. The linearity between passage segment number and $m$ has been shown in (huang2024homotopic). Figure 12 demonstrates that the linear increase of passage numbers with respect to $m$ also holds in the amended Gabriel condition. Because of the linearity between triangle and edge numbers in random Delaunay graphs, the linearity between Gabriel cells and passages is expected. The cell number is linear in the passage number and further, in the obstacle number in Figure 12. Its value is roughly half the passage number and close to the obstacle number.
@@
Statistics of passages in 3D environments are reported in Figure 12. Obstacles are assigned random heights in $[0,400]$. For comparison, planar passage and cell numbers on the base are listed. The data show that the 3D passage number is also linear in the obstacle number. In other words, $O(m)$ valid spatial passages exist. This linearity originates from the fact that all intermediate edges in the incremental Delaunay graph construction process are checked. The ratio of the spatial passage number to the base passage number is $1.7$ on average, suggesting good sparsity of spatial passages. Sparity of passages in planar and 3D space is crucial to enabling efficient planning in PTOPP.